# Semi-Supervised Clustering of Sparse Graphs: Crossing the Information-Theoretic Threshold

### 08 Jan 2024 v3

@@ Abstract
The stochastic block model is a canonical random graph model for clustering and community detection on network-structured data. Decades of extensive study on the problem have established many profound results, among which the phase transition at the Kesten-Stigum threshold is particularly interesting both from a mathematical and an applied standpoint. It states that no estimator based on the network topology can perform substantially better than chance on sparse graphs if the model parameter is below certain threshold. Nevertheless, if we slightly extend the horizon to the ubiquitous semi-supervised setting, such a fundamental limitation will disappear completely. We prove that with arbitrary fraction of the labels revealed, the detection problem is feasible throughout the parameter domain. Moreover, we introduce two efficient algorithms, one combinatorial and one based on optimization, to integrate label information with graph structures. Our work brings a new perspective to stochastic model of networks and semidefinite program research.
@@ 1 Introduction
Clustering has long been an essential subject of many research fields, such as machine learning, pattern recognition, data science, and artificial intelligence. In this section, we include some background information of its general setting and the semi-supervised approach.
@@ 1.1 Clustering on Graphs
The basic task of clustering or community detection in its general form is, given a (possibly weighted) graph, to partition its vertices into several densely connected groups with relatively weak external connectivity. This property is sometimes also called assortativity. Clustering and community detection are central problems in machine learning and data science with various applications in scientific research and industrial development. A considerable amount of data sets can be represented in the form of a network that consists of interacting nodes, and one of the first features of interest in such a situation is to understand which nodes are “similar”, as an end or as preliminary step towards other learning tasks. Clustering is used to find genetically similar sub-populations (Padhukasahasram, 2014), to segment images (Shi and Malik, 2000), to study sociological behavior (Newman et al., 2002), to improve recommendation systems (Linden et al., 2003), to help with natural language processing (Gopalan and Blei, 2013), etc. Since the 1970s, in different communities like social science, statistical physics and machine learning, a large diversity of algorithms have been developed such as:
@@
  • Optimization approaches identify the best cluster structures in regard to carefully designed cost functions, for instance, minimizing the cut (Hartuv and Shamir, 2000) and maximizing the Girvan-Newman modularity (Newman and Girvan, 2004).
Multiple lines of research intersect at a simple random graph model, which appears under many different names. In the machine learning and statistics literature around social networks, it is called the stochastic block model (SBM) (Holland et al., 1983), while it is known as the planted partition model (Bui et al., 1984) in theoretical computer science and referred to as inhomogeneous random graph model (Bollobás et al., 2007) in the mathematics literature. Moreover, it can also be interpreted as a spin-glass model (Decelle et al., 2011), a sparse-graph code (Abbe and Sandon, 2015) or a low-rank random matrix model (McSherry, 2001) and more.
@@
Note that if $p=q$, the planted bisection model is reduced to the so-called Erdős–Rényi random graph where all edges are generated independently with the same probability. Hence there exists no cluster structure. But if $p\gg q$, a typical graph will have two well-defined clusters. The scale of $p$ and $q$ also plays a significant role in the resulting graph, which will be discussed in detail later. They govern the amount of signal and noise in the graph generating process. As the key parameters that researchers work with, various regimes and thresholds are described by them.
The SBMs generate labels for vertices before the graph. The ground truth allows us to formally discuss the presence of community structures and measure the performance of algorithms in a meaningful way. It also supplies a natural basis to rigorously define the semi-supervised clustering problem. But as a parametrized statistical model, one can only hope that it serves as a good fit for the real data. Although not necessarily a realistic model, SBM provides us an insightful abstraction and captures some of the key phenomena (Mossel et al., 2015; Chen and Xu, 2016; Banks et al., 2016; Abbe et al., 2016; Abbe and Sandon, 2018).
Given a single realization of the graph $G$, our goal is to recover the labels $x$, up to certain level of accuracy. Formally, the ground truth of the underlying community structure is encoded using the vector $x\in\{+1,-1\}^{n}$, with $x_{i}=+1$ if $i\in S_{1}$, and $x_{i}=-1$ if $i\in S_{2}$. An estimator is a map $\hat{x}\mathrel{\mathop{:}}G_{n}\to\{+1,-1\}^{n}$ where $G_{n}$ is the space of graphs over $n$ vertices. We define the overlap between an estimator and the ground truth as
@@
In other words, exact recovery requires the entire partition to be correctly identified. Weak recovery only asks for substantially better performance than chance. In some literature, exact recovery is simply called recovery. Weak recovery is also called detection since as long as one can weakly recover the ground truth, there must exist community structure.
Note that if $G$ is an Erdős–Rényi random graph ($p=q$) then the overlap will be $o_{p}(1)$ for all estimators. This can be seen by noticing that $x$ and $G$ are independent in this setting and then applying Markov’s inequality. This has led to two additional natural questions about SBMs. On the one hand, we are interested in the distinguishability (or testing): is there a hypothesis test to distinguish random graph generated by Erdős–Rényi model (ERM) from random graph generated by SBMs that succeeds with high probability? One the other hand, we can ask about the model learnability (or parameter estimation): assuming that $G$ is drawn from an SBM ensemble, is it possible to obtain a consistent estimator for the parameters ($p,q$)? Although each of these questions is of independent interest, for symmetric SBMs with two symmetric communities (planted bisection model) the following holds (Abbe, 2018):
@@
The equivalence benefits the analysis of the model in turn. For example, direct analysis about weak recovery leads to the converse of phase transition theory (Mossel et al., 2015). While the achievability of the phase transition threshold (Massoulié, 2014) is proved by counting non-backtracking walks on the graph which gives consistent estimators of parameters. In the recent work (Montanari and Sen, 2016), hypothesis testing formulation is studied.
SBMs demonstrate the ‘fundamental limits’ of clustering and community detection as some necessary and sufficient conditions for feasibility of recovery, information-theoretically or computationally. Moreover, they are usually expressed in the form of phase transition. Sharp transitions exist in the parameter regimes between phases where the task is resolvable or not. For example, when the average degree grows as $\log n$, if the structure is sufficiently obvious then the underlying communities can be exactly recovered (Bickel and Chen, 2009), and the threshold at which this becomes possible has also been determined (Abbe et al., 2016). Above this threshold, efficient algorithms exist (Agarwal et al., 2017; Abbe and Sandon, 2015; Perry and Wein, 2017; Deng et al., 2021) that recover the communities exactly, labeling every vertex correctly with high probability; below this threshold, exact recovery is information-theoretically impossible. (Yun and Proutiere, 2016; Gao et al., 2017) studied a slightly weaker requirement that allows a vanishing proportion of misclassified nodes and proposed algorithms that achieve the optimal rates for various models. In contrast, we study a rather different regime where the reasonable question to ask is whether one can guarantee a misclassification proportion strictly less than $\frac{1}{2}$.
We acknowledge that SBMs in general are theoretical constructs which can not capture all the complicated situations emerging with real-life networks. But it succeeds in modeling the simplest situation where a community structure can be quantitatively described. In practice, instead of applying the model directly, it is advisable to initially conduct some heuristic and hierarchical analyses on the graph and see if the problem can be effectively reduced to those simple cases. For the graphs clearly violating the model assumptions, e.g., a sparse graph dominated by cliques, alternative and more tailored modeling approaches may be necessary.
@@ 1.2 Sparse Regime and Kesten-Stigum Threshold
Although it is mathematically challenging to work in the sparse regime, real-world data are likely to have bounded average degrees. (Leskovec et al., 2008) and (Strogatz, 2001) studied a large collection of the benchmark data sets, including power transmission networks, web link networks and complex biological systems, which had millions of nodes with average degree of no more than 20. For instance, the LinkedIn network they studied had approximately seven million nodes, but only 30 million edges.
The phase transition for weak recovery or detection in the sparse regime was first conjectured in the paper by Decelle, Krzakala, Moore, Zdeborová (Decelle et al., 2011), which sparked the modern study of clustering and SBMs. Their work is based on deep but non-rigorous insights from statistical physics, derived with the cavity method (a.k.a. belief propagation). Since then, extensive excellent research has been conducted to understand this fundamental limit, see e.g. (Mossel et al., 2015; Massoulié, 2014; Mossel et al., 2018; Abbe et al., 2020a). A key result is the following theorem.
@@
[Kesten-Stigum threshold] Let $\operatorname{\mathcal{G}}(n,a/n,b/n)$ be a symmetric SBM with two balanced clusters and $a,b=O(1)$. The weak recovery problem is solvable and efficiently so if and only if $(a-b)^{2}>2(a+b)$.
@@
$\text{SNR}\coloneqq(a-b)^{2}/[2(a+b)]$ (4)
It is worth noting that we only quoted the KS threshold for the two communities case ($k=2$). For sufficiently large $k$, namely $k\geq 5$, there is a ‘hard but detectable’ area where the weak recovery is information-theoretically possible, but computationally hard (Abbe and Sandon, 2018; Banks et al., 2016). This gap between the KS threshold and information-theoretic (IT) threshold only shows up in the constant degree regime, making it a fertile ground for studying the fundamental tradeoffs in community detection. We focus on the cardinal case, symmetric SBM with two balanced clusters, where two thresholds coincide and semi-supervised approach crosses them together.
@@
The terminology ‘KS threshold’ can be traced back to the work of Kesten and Stigum concerning reconstruction on infinite rooted trees in 1966 (Kesten and Stigum, 1966). The problem consists in broadcasting the root label of a tree with fixed degree $c$ down to its leaves, and trying to recover it from the leaves at large depth. We start with drawing the root label uniformly in $\{0,1\}$. Then, in a top-down manner, we independently label every child the same as its parent with probability $1-\epsilon$ and the opposite as its parent otherwise. Let $x^{(t)}$ denote the labels at depth $t$ in this tree with $t=0$ being the root. We say the reconstruction is solvable if $\lim_{t\to\infty}\operatorname{E}|\operatorname{E}(x^{(0)}|x^{(t)})-1/2|>0$ or, equivalently, $\lim_{t\to\infty}I(x^{(0)};x^{(t)})>0$, where $I$ is the mutual information. Although it was shown in the original paper, reconstruction is solvable when $c(1-2\epsilon)^{2}>1$, the non-reconstruction was proved 30 years later, namely it is not solvable if $c(1-2\epsilon)^{2}\leq 1$ (Bleher et al., 1995; Evans et al., 2000). Based on that finding, Mossel, Neeman, and Sly proved the converse part of Theorem 3 by coupling the local neighborhood of an SBM vertex with a Gaton-Watson tree with a Markov Process (Mossel et al., 2015). Inspired by this elegant approach, we propose our ‘census method’ to solve the semi-supervised clustering problem, and we will see in Section 3 how it works by amplifying revealed information with tree-like neighborhoods.
@@ 1.3 Basic Algorithms
Information-theoretic bounds can provide the impossibility side of phase transitions, but we still need specific efficient algorithms for the achievability side. One straight forward approach is spectral method. Under the Definition 1, let $A$ be the adjacency matrix of the graph $G\sim\operatorname{\mathcal{G}}(n,a/n,b/n)$, $a>b$. Up to reordering indices, its expectation is a block matrix except for the diagonal,
$\operatorname{E}A\approx\frac{1}{n}\begin{pmatrix}a&b\\ b&a\end{pmatrix}\otimes I_{n/2\times n/2}$ (5)
which has three eigenvalues, $(a+b)/n>(a-b)/n>0$. 00 has multiplicity $n-2$ and the eigenvector associated with the second largest eigenvalue is $\left(\begin{smallmatrix}\mathbf{1}_{n/2}\\ -\mathbf{1}_{n/2}\end{smallmatrix}\right)$ which is consistent with the ground truth of the labels. But we do not observe the expected adjacency matrix. Instead, we only have access to one realization of the model. In modern terms, community detection is a ‘one-shot learning’ task. But one can still hope that $A-\operatorname{E}A$ is small and the second eigenvector of $A$ gives a reasonable estimator. For example, denoting the ordered eigenvalues of $\operatorname{E}A$ and $A$ as $\{\lambda_{i}\}$’s and $\{\hat{\lambda}_{i}\}$’s respectively, the Courant-Fischer-Weyl min-max principle implies
@@
Recall that the operator norm of a symmetric matrix $M$, with $\xi_{i}(M)$ being its $i$-th largest eigenvalue, is $\|M\|_{\text{op}}=\max(\xi_{1}(M),-\xi_{n}(M))$. If one can bound $\|A-\operatorname{E}A\|_{\text{op}}$ by half of the least gap between the three eigenvalues mentioned above, the order will be preserved. Then the Davis-Kahan theorem guarantees the eigenvectors are correlated. Namely, if $\theta$ denotes the angle between the second eigenvectors (spectral estimator and ground truth), we have
@@
This nice and simple approach stops working as we step into the sparse regime (Feige and Ofek, 2005; Coja-Oghlan, 2009; Keshavan et al., 2009; Decelle et al., 2011; Krzakala et al., 2013). The main reason is that leading eigenvalues of $A$ are about the order of square root of the maximum degree. High degree vertices mess up the desired order of eigenvalues. In particular, for Erdős–Rényi random graphs ($\operatorname{\mathcal{G}}(n,d/n)$), we have $\hat{\lambda}_{1}=(1+o(1))\sqrt{\log n/\log\log n}$ almost surely (Krivelevich and Sudakov, 2003). Furthermore, the leading eigenvectors are concentrated at these high degree ‘outliers’ and contain no structural information of the underlying model.
Take the star graph for example, where we assume that only the first node is connected to $k$ neighbors. It is easy to see that the corresponding adjacency matrix has eigenvalue $\sqrt{k}$ and eigenvector $(\sqrt{k},1,\dots,1)$. Various interesting spectrum based methods are proposed to overcome this challenge (Mossel et al., 2018; Massoulié, 2014; Bordenave et al., 2015). The key idea is to replace adjacency with some combinatorically constructed matrices. However, they typically rely on model statistics and underlying probabilistic assumptions, which leads to the problem of adversarial robustness. For example, they are non-robust to ’helpful’ perturbations. Namely, if we let an adversary to perform following changes on the graph: (1) adding edges within communities and/or (2) removing edges across communities, spectral approaches are going to fail. It is surprising since, intuitively, these changes help to emphasize community structures.
Meanwhile, semidefinite programming (SDP) sheds the light on how we may be able to overcome the limitations of spectral algorithms, which is shown to be robust when SNR is sufficiently large (Moitra et al., 2016). In fact, it is another major line of work on clustering and community detection concerning performance of SDPs on SBMs. While a clear picture of the unbounded degree case is figured out in (Abbe et al., 2016; Hajek et al., 2016; Amini and Levina, 2018; Bandeira, 2018; Agarwal et al., 2017; Perry and Wein, 2017), the results for sparse networks are more complicated. (Guédon and Vershynin, 2014) proved a sub-optimal condition, SNR $\geq 10^{4}$, using Grothendieck inequality. Then, with a Lindeberg interpolation process (Tao, 2011), Montanari et al. proved that a SDP algorithm as proposed in (Montanari and Sen, 2016) is nearly optimal for large bounded average degree by transferring analysis of the original SDPs to the analysis of SDPs of Gaussian random matrices.
@@
The necessary degree $d^{*}$ depends on the choice of $\epsilon$ and goes to infinity as $\epsilon\to 0$. If we fix $d$ and view $\epsilon$ as its function, the condition becomes $\text{SNR}\geq 1+o_{d}(1)$. Numerical estimation and non-rigorous statistical mechanism approximation suggest that it is at most $2\%$ sub-optimal. This result seems to be the ceiling of SDP according to the preliminary calculation in (Javanmard et al., 2016). Moreover, they address the irregularity of high degree nodes by showing SDPs return similar results for Erdős–Rényi random graphs and random regular graphs, which appear to be sensitive only to the average degree. See Section 4 for more discussion on the estimation. Inspired by their work, we propose a natural modification of SDP to incorporate revealed labels in the semi-supervised setting and show that it not only achieves, but even crosses, the KS threshold. In turn, our result brings a new perspective to study the (non-)achievability and robustness of (unsupervised) SDPs.

### 28 Feb 2024 v4

@@ Abstract
The stochastic block model is a canonical random graph model for clustering and community detection on network-structured data. Decades of extensive study on the problem have established many profound results, among which the phase transition at the Kesten-Stigum threshold is particularly interesting both from a mathematical and an applied standpoint. It states that no estimator based on the network topology can perform substantially better than chance on sparse graphs if the model parameter is below a certain threshold. Nevertheless, if we slightly extend the horizon to the ubiquitous semi-supervised setting, such a fundamental limitation will disappear completely. We prove that with an arbitrary fraction of the labels revealed, the detection problem is feasible throughout the parameter domain. Moreover, we introduce two efficient algorithms, one combinatorial and one based on optimization, to integrate label information with graph structures. Our work brings a new perspective to the stochastic model of networks and semidefinite program research.
@@ 1 Introduction
Clustering has long been an essential subject of many research fields, such as machine learning, pattern recognition, data science, and artificial intelligence. In this section, we include some background information on its general setting and the semi-supervised approach.
@@ 1.1 Clustering on Graphs
The basic task of clustering or community detection in its general form is, given a (possibly weighted) graph, to partition its vertices into several densely connected groups with relatively weak external connectivity. This property is sometimes also called assortativity. Clustering and community detection are central problems in machine learning and data science with various applications in scientific research and industrial development. A considerable amount of data sets can be represented in the form of a network that consists of interacting nodes, and one of the first features of interest in such a situation is to understand which nodes are “similar”, as an end or as a preliminary step towards other learning tasks. Clustering is used to find genetically similar sub-populations (Padhukasahasram, 2014), to segment images (Shi and Malik, 2000), to study sociological behavior (Newman et al., 2002), to improve recommendation systems (Linden et al., 2003), to help with natural language processing (Gopalan and Blei, 2013), etc. Since the 1970s, in different communities like social science, statistical physics, and machine learning, a large diversity of algorithms have been developed such as:
@@
  • Optimization approaches identify the best cluster structures regarding carefully designed cost functions, for instance, minimizing the cut (Hartuv and Shamir, 2000) and maximizing the Girvan-Newman modularity (Newman and Girvan, 2004).
Multiple lines of research intersect at a simple random graph model, which appears under many different names. In the machine learning and statistics literature around social networks, it is called stochastic block model (SBM) (Holland et al., 1983), while it is known as the planted partition model (Bui et al., 1984) in theoretical computer science and referred to as inhomogeneous random graph model (Bollobás et al., 2007) in the mathematics literature. Moreover, it can also be interpreted as a spin-glass model (Decelle et al., 2011), a sparse-graph code (Abbe and Sandon, 2015), a low-rank random matrix model (McSherry, 2001), and more.
@@
Note that if $p=q$, the planted bisection model is reduced to the so-called Erdős–Rényi random graph where all edges are generated independently with the same probability. Hence there exists no cluster structure. But if $p\gg q$, a typical graph will have two well-defined clusters. The scale of $p$ and $q$ also plays a significant role in the resulting graph, which will be discussed in detail later. They govern the amount of signal and noise in the graph’s generating process. As the key parameters that researchers work with, they depict various regimes and thresholds.
The SBMs generate labels for vertices before the graph. The ground truth allows us to formally discuss the presence of community structures and measure the performance of algorithms in a meaningful way. It also supplies a natural basis to rigorously define the semi-supervised clustering problem. But as a parametrized statistical model, one can only hope that it serves as a good fit for the real data. Although not necessarily a realistic model, SBM provides us with an insightful abstraction and captures some of the key phenomena (Mossel et al., 2015; Chen and Xu, 2016; Banks et al., 2016; Abbe et al., 2016; Abbe and Sandon, 2018).
Given a single realization of a graph $G$, our goal is to recover the labels $x$, up to a certain level of accuracy. Formally, the ground truth of the underlying community structure is encoded using the vector $x\in\{+1,-1\}^{n}$, with $x_{i}=+1$ if $i\in S_{1}$, and $x_{i}=-1$ if $i\in S_{2}$. An estimator is a map $\hat{x}\mathrel{\mathop{:}}G_{n}\to\{+1,-1\}^{n}$ where $G_{n}$ is the space of graphs over $n$ vertices. We define the Overlap between an estimator and the ground truth as
@@
In other words, exact recovery requires the entire partition to be correctly identified. Weak recovery only asks for substantially better performance than chance. In some literature, exact recovery is simply called recovery. Weak recovery is also called detection since as long as one can weakly recover the ground truth, there must exist a community structure.
Note that if $G$ is an Erdős–Rényi random graph ($p=q$) then the overlap will be $o_{p}(1)$ for all estimators. This can be seen by noticing that $x$ and $G$ are independent in this setting and then applying Markov’s inequality. This has led to two additional natural questions about SBMs. On the one hand, we are interested in the distinguishability (or testing): is there a hypothesis test to distinguish a random graph generated by the Erdős–Rényi model (ERM) from a random graph generated by the SBM, which succeeds with high probability? On the other hand, we can ask about the model learnability (or parameter estimation): assuming that $G$ is drawn from an SBM ensemble, is it possible to obtain a consistent estimator for the parameters ($p,q$)? Although each of these questions is of independent interest, for symmetric SBMs with two symmetric communities (planted bisection model) the following holds (Abbe, 2018):
@@
Such equivalence benefits our understanding of the model in turn. For example, direct analysis of weak recovery leads to the converse of phase transition theory (Mossel et al., 2015). The achievability of the phase transition threshold (Massoulié, 2014) is proved by counting non-backtracking walks on the graph which gives consistent estimators of parameters. In the recent work (Montanari and Sen, 2016), hypothesis testing formulation is studied.
SBMs demonstrate the ‘fundamental limits’ of clustering and community detection as some necessary and sufficient conditions for the feasibility of recovery, information-theoretically or computationally. Moreover, they are usually expressed in the form of phase transition. Sharp transitions exist in the parameter regimes between phases where the task is resolvable or not. For example, when the average degree grows as $\log n$, if the structure is sufficiently obvious then the underlying communities can be exactly recovered (Bickel and Chen, 2009), and the threshold at which this becomes possible has also been determined (Abbe et al., 2016). Above this threshold, efficient algorithms exist (Agarwal et al., 2017; Abbe and Sandon, 2015; Perry and Wein, 2017; Deng et al., 2021) that recover the communities exactly, labeling every vertex correctly with high probability; below this threshold, exact recovery is information-theoretically impossible. (Yun and Proutiere, 2016; Gao et al., 2017) studied a slightly weaker requirement that allows a vanishing proportion of misclassified nodes and proposed algorithms that achieve the optimal rates for various models. In contrast, we study a rather different regime where the reasonable question to ask is whether one can guarantee a misclassification proportion strictly less than $\frac{1}{2}$.
We acknowledge that SBMs in general are theoretical constructs that can not capture all the complicated situations emerging with real-life networks. However, it succeeds in modeling the simplest situation where a community structure can be quantitatively described. In practice, instead of applying the model directly, it is advisable to initially conduct some heuristic and hierarchical analyses on the graph and see if the problem can be effectively reduced to those simple cases. For the graphs violating the model assumptions, e.g., a sparse graph dominated by cliques, alternative and tailored modeling may be necessary.
@@ 1.2 Sparse Regime and Kesten-Stigum Threshold
Although it is mathematically challenging to work in the sparse regime, real-world data are likely to have bounded average degrees. (Leskovec et al., 2008) and (Strogatz, 2001) studied a large collection of benchmark data sets, including power transmission networks, website link networks, and complex biological systems, which had millions of nodes with an average degree of no more than 20. For instance, the LinkedIn network they studied had approximately seven million nodes, but only 30 million edges.
The phase transition for weak recovery or detection in the sparse regime was first conjectured in the paper by Decelle, Krzakala, Moore, Zdeborová (Decelle et al., 2011), which sparked the modern study of clustering and SBMs. Their work is based on deep but non-rigorous insights from statistical physics, derived with the cavity method (a.k.a. belief propagation). Since then, extensive excellent research has been conducted to understand this fundamental limit, e.g., (Mossel et al., 2015; Massoulié, 2014; Mossel et al., 2018; Abbe et al., 2020a). A key result is the following theorem.
@@
[Kesten-Stigum threshold] Let $\operatorname{\mathcal{G}}(n,a/n,b/n)$ be a symmetric SBM with two balanced clusters and $a,b=O(1)$. The weak recovery problem is solvable and efficiently so, if and only if $(a-b)^{2}>2(a+b)$.
@@
$\text{SNR}\coloneqq(a-b)^{2}/[2(a+b)].$ (4)
It is worth noting that we only quoted the KS threshold for the two communities case ($k=2$). For sufficiently large $k$, namely $k\geq 5$, there is a ‘hard but detectable’ area where the weak recovery is information-theoretically possible, but computationally hard (Abbe and Sandon, 2018; Banks et al., 2016). This gap between the KS threshold and the information-theoretic (IT) threshold only shows up in the constant degree regime, making it a fertile ground for studying the fundamental tradeoffs in community detection. We focus on the cardinal case, symmetric SBM with two balanced clusters, where two thresholds coincide and a semi-supervised approach crosses them together.
@@
The terminology ‘KS threshold’ can be traced back to the work of Kesten and Stigum concerning the reconstruction of infinite rooted trees in 1966 (Kesten and Stigum, 1966). The problem consists of broadcasting the root label of a tree with a fixed degree $c$ down to its leaves and trying to recover it from the leaves at a large depth. We start with drawing the root label uniformly in $\{0,1\}$. Then, in a top-down manner, we independently label every child the same as its parent with probability $1-\epsilon$ and the opposite as its parent otherwise. Let $x^{(t)}$ denote the labels at depth $t$ in this tree with $t=0$ being the root. We say the reconstruction is solvable if $\lim_{t\to\infty}\operatorname{E}|\operatorname{E}(x^{(0)}|x^{(t)})-1/2|>0$ or, equivalently, $\lim_{t\to\infty}I(x^{(0)};x^{(t)})>0$, where $I$ is the mutual information. Although it was shown in the original paper, reconstruction is solvable when $c(1-2\epsilon)^{2}>1$, the non-reconstruction was proved 30 years later, namely, it is not solvable if $c(1-2\epsilon)^{2}\leq 1$ (Bleher et al., 1995; Evans et al., 2000). Based on that finding, Mossel, Neeman, and Sly proved the converse part of Theorem 3 by coupling the local neighborhood of an SBM vertex with a Gaton-Watson tree with a Markov Process (Mossel et al., 2015). Inspired by this elegant approach, we propose our ‘census method’ to solve the semi-supervised clustering problem, and we will see in Section 3 how it works by amplifying revealed information with tree-like neighborhoods.
@@ 1.3 Basic Algorithms
Information-theoretic bounds can provide the impossibility side of phase transitions, but we still need specific efficient algorithms for the achievability side. One straightforward approach is the spectral method. Under Definition 1, let $A$ be the adjacency matrix of the graph $G\sim\operatorname{\mathcal{G}}(n,a/n,b/n)$, $a>b$. Up to reordering indices, its expectation is a block matrix except for the diagonal,
$\operatorname{E}A\approx\frac{1}{n}\begin{pmatrix}a&b\\ b&a\end{pmatrix}\otimes I_{n/2\times n/2},$ (5)
which has three eigenvalues, $(a+b)/n>(a-b)/n>0$. 00 has multiplicity $n-2$ and the eigenvector associated with the second largest eigenvalue is $\left(\begin{smallmatrix}\mathbf{1}_{n/2}\\ -\mathbf{1}_{n/2}\end{smallmatrix}\right)$ which is consistent with the ground truth of the labels. However, we do not observe the expected adjacency matrix. Instead, we only have access to one realization of the model. In modern terms, community detection is a ‘one-shot learning’ task. But one can still hope that $A-\operatorname{E}A$ is small and the second eigenvector of $A$ gives a reasonable estimator. For example, denoting the ordered eigenvalues of $\operatorname{E}A$ and $A$ as $\{\lambda_{i}\}$’s and $\{\hat{\lambda}_{i}\}$’s respectively, the Courant-Fischer-Weyl min-max principle implies
@@
Recall that the operator norm of a symmetric matrix $M$ is $\|M\|_{\text{op}}=\max(\xi_{1}(M),-\xi_{n}(M))$ where $\xi_{i}(M)$ denotes the $i$-th largest eigenvalue of $M$. If one can bound $\|A-\operatorname{E}A\|_{\text{op}}$ by half of the least gap between the three eigenvalues mentioned above, the order will be preserved. Then the Davis-Kahan theorem guarantees the eigenvectors are correlated. Namely, if $\theta$ denotes the angle between the second eigenvectors (spectral estimator and ground truth), we have
@@
This nice and simple approach stops working as we step into the sparse regime (Feige and Ofek, 2005; Coja-Oghlan, 2009; Keshavan et al., 2009; Decelle et al., 2011; Krzakala et al., 2013). The main reason is that leading eigenvalues of $A$ are about the order of the square root of the maximum degree. High-degree vertices mess up the desired order of eigenvalues. In particular, for Erdős–Rényi random graphs ($\operatorname{\mathcal{G}}(n,d/n)$), we have $\hat{\lambda}_{1}=(1+o(1))\sqrt{\log n/\log\log n}$ almost surely (Krivelevich and Sudakov, 2003). Furthermore, the leading eigenvectors are concentrated at these ‘outliers’ with high degree and contain no structural information of the underlying model.
Take the star graph for example, where we assume that only the first node is connected to $k$ neighbors. It is easy to see that the corresponding adjacency matrix has eigenvalue $\sqrt{k}$ and eigenvector $(\sqrt{k},1,\dots,1)$. Various interesting spectrum-based methods are proposed to overcome this challenge (Mossel et al., 2018; Massoulié, 2014; Bordenave et al., 2015). The key idea is to replace adjacency with some combinatorically constructed matrices. However, they typically rely on model statistics and underlying probabilistic assumptions, which leads to the problem of adversarial robustness. For example, they are non-robust to ’helpful’ perturbations. Namely, if we allow an adversary to perform the following changes on the graph: (1) adding edges within communities and/or (2) removing edges across communities, spectral approaches are going to fail. It is surprising since, intuitively, these changes help to emphasize community structures.
Meanwhile, semidefinite programming (SDP) sheds light on how we may be able to overcome the limitations of spectral algorithms, which are shown to be robust when SNR is sufficiently large (Moitra et al., 2016). It is another major line of work on clustering and community detection concerning the performance of SDPs on SBMs. While a clear picture of the unbounded degree case has been figured out by (Abbe et al., 2016; Hajek et al., 2016; Amini and Levina, 2018; Bandeira, 2018; Agarwal et al., 2017; Perry and Wein, 2017), the results for sparse networks are more complicated. (Guédon and Vershynin, 2014) proved a sub-optimal condition, SNR $\geq 10^{4}$, using Grothendieck inequality. Then, with a Lindeberg interpolation process (Tao, 2011), Montanari et al. proved that an SDP algorithm as proposed by (Montanari and Sen, 2016) is nearly optimal for the case of large bounded average degree by transferring analysis of the original SDPs to the analysis of SDPs of Gaussian random matrices.
@@
The necessary degree $d^{*}$ depends on the choice of $\epsilon$ and goes to infinity as $\epsilon\to 0$. If we fix $d$ and view $\epsilon$ as its function, the condition becomes $\text{SNR}\geq 1+o_{d}(1)$. Numerical estimation and non-rigorous statistical mechanism approximation suggest that it is at most $2\%$ sub-optimal. This result seems to be the ceiling of SDP according to the preliminary calculation from (Javanmard et al., 2016). Moreover, they address the irregularity of high-degree nodes by showing SDPs return similar results for Erdős–Rényi random graphs and random regular graphs, which appear to be sensitive only to the average degree. See Section 4 for more discussion on the estimation. Inspired by their work, we propose a natural modification of SDP to incorporate revealed labels in the semi-supervised setting and show that it not only achieves but even crosses, the KS threshold. In turn, our result brings a new perspective to study the (non-)achievability and robustness of (unsupervised) SDPs.